OpenStudy (anonymous):

How to solve: 4x^2/12x^9 Please run me through the steps so I can learn to do it myself. Thanks.

OpenStudy (anonymous):



OpenStudy (lgbasallote):

pull out the constant first... 4/12(x^2/x^9) simplify the constant... 1/3(x^2/x^9) perform the operation of the variables 1/3(x^2/x^9) = 1/3(x^2-9) = 1/3(x^-7) simplify... 1/(3x^7)

OpenStudy (anonymous):

Thank you!